the partial pressure of CH4(g) is 0.175 atm and that of O2(g) is 0.250 atm in a mixture of the two gases.
A. what is the mole fraction of each gas in the mixture ?
B. If the mixture occupies a volume of 10.5 L at 65°C ,calculate the total number of moles of gas in the mixture
C. Calculate the number of grams of each gas in the mixture.
(answer in print not cursive)
Answer #1

(A.) Partial pressure CH4 = 0.175 atm

Partial pressure O2 = 0.250 atm

Total pressure = (Partial pressure CH4) + (Partial pressure O2)

Total pressure = (0.175 atm) + (0.250 atm)

Total pressure = 0.425 atm

mole fraction CH4 = (Partial pressure CH4) / (Total pressure)

mole fraction CH4 = (0.175 atm) / (0.425 atm)

mole fraction CH4 = 0.412

mole fraction O2 = 1 - (mole fraction CH4)

mole fraction O2 = 1 - (0.412)

mole fraction O2 = 0.588

(B.) Volume of mixture = 10.5 L

Temperature of mixture = 65oC = (273 + 65) K = 338 K

According to ideal gas law,

Total moles = [(Total pressure) * (volume of mixture)] / [(R) * (absolute temperature of mixture)]

where R = gas constant = 0.0821 L-atm/mol-K

Total moles = [(0.425 atm) * (10.5 L)] / [(0.0821 L-atm/mol-K) * (338 K)]

Total moles = 0.161 mol

(C.) moles CH4 in mixture = (mole fraction CH4) * (total moles)

moles CH4 in mixture = (0.412) * (0.161 mol)

moles CH4 in mixture = 0.0663 mol

mass CH4 in mixture = (moles CH4 in mixture) * (molar mass CH4)

mass CH4 in mixture = (0.0663 mol) * (16.0 g/mol)

mass CH4 in mixture = 1.06 g

moles O2 in mixture = (mole fraction O2) * (total moles)

moles O2 in mixture = (0.588) * (0.161 mol)

moles O2 in mixture = 0.0947 mol

mass O2 in mixture = (moles O2 in mixture) * (molar mass O2)

mass O2 in mixture = (0.0947 mol) * (32.0 g/mol)

mass O2 in mixture = 3.03 g